OpenStudy (anonymous):

what is the solution to this system y = x² + 3x – 4 y = 2x + 2

OpenStudy (anonymous):

what is the solution to this system y = x² + 3x – 4 = Y=x^2+3x-4

OpenStudy (anonymous):

y = 2x + 2 =y=2(x+1)

hero (hero):

Set y = y x^2 + 3x - 4 = 2x + 2 Then solve for x

OpenStudy (anonymous):

how do I solve for x?

hero (hero):

subtract 2x+2 from both sides to get x^2 + x - 2 = 0 Then factor by grouping: x^2 + 2x - x - 2 = 0 x(x+2)-1(x+2) = 0 (x+2)(x-1) = 0 x = -2,1